# The Distributional Uncertainty of the SHAP score in Explainable Machine Learning

### v1

@@ 4.3 Related problems
[…]
@@
The proof follows the same techniques as the proof for Theorem 7, through a reduction from VERTEX-COVER. For the case of REGION-IRRELEVANCY we device a model $M$ such that
@@ 5 Experimentation
As a case study, we are going to use the California Housing dataset Nugent, C. (2018), a comprehensive collection of housing data within the state of California containing 20,640 samples and eight features. Our choice of dataset relies mainly on the fact that this dataset has already been considered in the context of SHAP score computation Bertossi and León (2023) and its size and number of features allow us to compute most of the proposed parameters (as the SHAP polynomial itself, for each feature) in a reasonable time. Nonetheless, we recall that the proposed framework can be applied to any dataset, as long as there is some uncertainty on the real distribution of feature values and uncertainty regions can be estimated for it (e.g. via sampling the data as we do here).
Note that, while there is no reason to expect that the probabilities of each feature are independent of each other in the California Housing dataset, we make this assumption in our framework (which is a common one in the literature Štrumbelj and Kononenko (2014); Shrikumar et al. (2017)).
@@ 5.1 Objectives
The purpose of this experiment is to use a real dataset to simulate a situation where we have uncertainty over the proportions of each feature in the dataset. We want to derive suitable hyperrectangles representing the distribution uncertainty where our extended concepts of SHAP scores apply, and compare these scores against the usual, point-like SHAP score, in order to reveal cases where these new hypervolume scores are more informative than the traditional scheme. We expect our proposal to be able to detect features whose ranking is vulnerable to small distribution shifts, and aim to study how such sensitivity starts to vanish as we reduce the uncertainty over the distribution (i.e. as the hyperrectangles get smaller).
@@ 5.2 Methodology
We pick 20 different entities uniformly at random and compute the SHAP score for each of these entities and for each feature at all the vertices of the different hyperrectangles. Instead of using the polynomial-time algorithm for the SHAP value computation over decision trees, we actually compute an algebraic expression for the SHAP polynomial for each pair of entity and feature, simplify it, and then use it to compute any desired SHAP score given any product distribution. [^7] The complete code for the experiments can be found at:
https://git.exactas.uba.ar/scifuentes/fuzzyshapley.
@@ 5.3 Results
We can also observe in Figure 1 that something similar happens as well for feature housing_median_age. When $p=0.4\%$ its SHAP interval intersects with the intervals of the three features longitude, latitude and population. Nonetheless, by inspecting the difference of the scores at the vertices of $\mathcal{I}$ it can be seen that there is no point in which housing_median_age is ranked below the other features (i.e. housing_median_age dominates all three features). Meanwhile, if we compare two of the other features such as longitude and latitude, we observe that they have a relevant intersection even when $p=6.4\%$, and in 179 out of the 256 vertices latitude is ranked below longitude.
@@ 6 Conclusions
We have analyzed how SHAP scores for classification problems depend on the underlying distribution over the entity population when the distribution varies over a given uncertainty region. As a first and important step, we focused on product distributions and hyperrectangles as uncertainty regions, and obtained algorithmic and complexity results for fundamental problems related to how SHAP scores vary under these conditions. We showed through experimentation that considering uncertainty regions has an impact on feature rankings for a non-negligible percentage of the entities, and that the solutions to our proposed problems can provide insight on the relative rankings even in the presence of uncertainty.
There are several problems left open by our work. First, going beyond the case of product distributions is a natural next step, and similarly for non-binary features and labels. Second, It would be interesting to know how a local perturbation of a given distribution in the uncertainty region affects the SHAP scores, or their rankings. This would be a proper robustness analysis with respect to the distribution (as opposed to how SHAP scores vary in a region). Of course, this is a different problem from the most common one of robustness with respect to the perturbation of an input entity. Lastly, it would be interesting to analyze others proposed scores (e.g. LIME Ribeiro et al. (2016), RESP Bertossi et al. (2020), Kernel-SHAP Lundberg and Lee (2017)) in the setting of distributional uncertainty.
@@ Acknowledgments
This research has been supported by an STIC-AmSud program, project AMSUD210006. We appreciate the code made available by Jorge E. León (UAI, Chile) for SHAP computation with the same dataset used in our work. L. Bertossi has been supported by NSERC-DG 2023-04650 and the Millennium Institute of Foundational Research on Data (IMFD, Chile). Romero is funded by the National Center for Artificial Intelligence CENIA FB210017, BasalANID. Pardal is funded by DFG VI 1045-1/1. Abriola, Cifuentes, Martinez, and Pardal are funded by FONCyT, ANPCyT and CONICET, in the context of the projects PICT 01481 and PIP 11220200101408CO.
@@ Appendix A Appendix
Although the proof of this proposition can be found in Laneve et al. [2010], we show a different argument that we consider simpler.
@@
Theorem 5. Given a multilinear polynomial $f$, a hyperrectangle =[$\mathcal{I}={}_{i=1}^{n}[a_{i},b_{i}]$, and a rational $b$, the problem of deciding whether there is an $\textbf{x}\in\mathcal{I}$ such that $f(\textbf{x})\geq b$ is NP-complete.
@@
The problem is in NP, and the proof of this fact was already sketched in Section 4.1: due to Proposition 4, if $f$ attains a value larger than $b$, then there is a vertex x from $\mathcal{I}$ such that $f(\textbf{x})\geq b$. Therefore, a non-deterministic Turing Machine can guess which vertex it is and perform the evaluation.
@@
We prove the hardness of REGION-MAX-SHAP through a reduction from VERTEX-COVER. Let $G=(V,E)$ be a graph and $k\geq 1$ be an integer. We shall define a classifier $M$, an entity $e_{0}$, a feature $x_{0}$, an interval $\mathcal{I}$, and a bound $b$ such that $G$ has a vertex cover of size at most $k$ if and only if there is a $\textbf{p}\in\mathcal{I}$ such that $Shap_{M,e_{0},x_{0}}(\textbf{p})\geq b$.
The set of features is $X\vcentcolon=\{x_{0}\}\cup V\cup\{w\}$ where $V$ is the set of nodes of $G$ and $n\vcentcolon=|V|$. We consider the null entity $e_{0}$ that assigns 00 to all features, and the feature $x_{0}$ to define the SHAP polynomial. The hyperrectangle is $\mathcal{I}=[1,1]\times{}_{i=1}^{n}[0,1]\times[\varepsilon,\varepsilon]$. The value of $\varepsilon$ and the bound $b$ will be defined later. Recall that we can assume that the maximum of $Shap_{M,e_{0},x_{0}}$ is attained at a vertex of $\mathcal{I}$, i.e. at a point where $p_{x_{0}}=1$, $p_{w}=\varepsilon$ and $p_{v}\in\{0,1\}$ for all $v\in V$. The probabilities $(p_{v})_{v\in V}$ will be used to “mark” which nodes belong to the vertex cover. Formally, for a subset of nodes $C\subseteq V$ of $G$, we associate the vector $\textbf{p}^{C}=(p_{x})_{x\in X}$ such that $p_{x_{0}}=1$, $p_{w}=\varepsilon$, and for $v\in V$, we have $p_{v}=0$ if $v\in C$, and $p_{v}=1$ otherwise. Note that this gives us a bijection between subsets of nodes of $G$ and vertices of $\mathcal{I}$.
Intuitively, we device the classifier $M$ such that $Shap_{M,e_{0},x_{0}}$ has the form
@@
where $\ell$ is the size of the subset $C$ and the term $T_{n,\ell}$ grows as $\ell$ grows. The terms $I_{u,v}$ can be made sufficiently large by choosing $\varepsilon$ properly. We choose the bound $b$ to be $T_{n,k}$. If $C$ is a vertex cover, then each term $p_{u}p_{v}I_{u,v}=0$ and hence the first sum of $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})$ is 00. On the other hand, if $C$ is not a vertex cover, we get a small negative value for $Shap_{M,e_{0},x_{0}}$, smaller than $b$ in particular, as we get penalized by at least one $I_{u,v}$ (from an uncovered edge $\{u,v\}$). Hence the only way to obtain $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})\geq b$ is to pick $C$ to be, in the first place, a vertex cover, and secondly, one of size $\ell\leq k$.
@@
Recall $\ell$ denotes the size of $C$. We pick $\varepsilon\vcentcolon=\frac{c_{n-1}}{\sum_{i=0}^{k}\binom{k}{i}c_{n-i}}$ and $b\vcentcolon=-\varepsilon\sum_{i=0}^{k}\binom{k}{i}c_{n-i}=-c_{n-1}$. These values can be computed in $poly(n)$ because $k\leq n$. Similarly, the rest of the reduction can be performed in polynomial time.
Now we show the correctness of the reduction: $G$ has a vertex cover of size at most $k$ if and only if there is a $\textbf{p}\in\mathcal{I}$ such that $Shap_{M,e_{0},x_{0}}(\textbf{p})\geq b$. Suppose $G$ has a vertex cover of size at most $k$. Then it has a vertex cover $C$ of size precisely $k$. We can take the point $\textbf{p}^{C}$ since:
@@
$=b.$
Suppose now that there is no vertex cover of $G$ of size at most $k$. We claim that $Shap_{M,e_{0},x_{0}}(\textbf{p})<b$ for all $\textbf{p}\in\mathcal{I}$. It suffices to check this for the vertices of $\mathcal{I}$, that is, for all vectors of the form $\textbf{p}^{C}$ for subsets of nodes $C\subseteq V$. Now, take an arbitrary subset $C$. Suppose first that $C$ is not a vertex cover, and that the edge $\{u,v\}$ is not covered, then:
@@
Note that $I_{u,v}\geq c_{n-1}$ as $c_{n-1}$ participates in the sum when $S=(V\setminus\{u,v\})\cup\{w\}$. It follows that $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})<-c_{n-1}=b$.
@@
$Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})<-\varepsilon\sum_{i=0}^{k}\binom{k}{i}c_{n-i}=b.$
@@ A.3 The problem REGION-MIN-SHAP
Thus, $b$ is an upper bound for $Shap_{M,e,x}(\textbf{p})$ for $\textbf{p}\in\mathcal{I}$ if and only if $-b$ is a lower bound of $Shap_{1-M,e,x}(\textbf{p})$ over the same hyperrectangle. ∎
@@
The idea behind our construction is similar to the one in the proof of Theorem 7. In this case, we device a model $M$ such that
@@
Finally, let $\mathcal{D}(S)=\phi_{M^{\prime},e^{\prime}}(S\cup\{x\})-\phi_{M^{\prime},e^{\prime}}(S)=\mathcal{D}(S\setminus\{w\})$ and observe that
@@
$=\sum_{\begin{subarray}{c}S\subseteq X^{\prime}\setminus\{x,w\}\end{subarray}}c_{|S|+1,n+1}\mathcal{D}(S)+c_{|S|,n+1}\mathcal{D}(S)$ (5)

### v2

@@ 4.3 Related problems
Let $\cal F$ be a class of classifiers for which computing the SHAP score for given product distributions can be done in polynomial time. Then the problems REGION-AMBIGUITY and REGION-IRRELEVANCY are in NP for the class $\mathcal{F}$, while the FEATURE-DOMINANCE is in coNP [^7] The proof for FEATURE-DOMINANCE follows by observing that $\textit{diff}(\textbf{p})=Shap_{M,e,x}(\textbf{p})-Shap_{M,e,y}(\textbf{p})$ is again a multilinear polynomial..
@@
The proof follows the same techniques as the proof for Theorem 5, through a reduction from VERTEX-COVER. For the case of REGION-IRRELEVANCY we devise a model $M$ such that
@@ 5 Experimentation
As a case study, we are going to use the California Housing dataset [19], a comprehensive collection of housing data within the state of California containing 20,640 samples and eight features [^9] The code developed for the experimentation can be found at https://git.exactas.uba.ar/scifuentes/fuzzyshapley.. Our choice of dataset relies mainly on the fact that this dataset has already been considered in the context of SHAP score computation [6] and its size and number of features allow us to compute most of the proposed parameters (as the SHAP polynomial itself, for each feature) in a reasonable time. Nonetheless, we recall that the proposed framework can be applied to any dataset, as long as there is some uncertainty on the real distribution of feature values and uncertainty regions can be estimated for it (e.g., via sampling the data as we do here).
Note that, while there is no reason to expect that the probabilities of each feature are independent of each other in the California Housing dataset, we make this assumption in our framework (which is a common one in the literature [26, 24]).
@@ 5.1 Objectives
The purpose of this experiment is to use a real dataset to simulate a situation where we have uncertainty over the proportions of each feature in the dataset. We want to derive suitable hyperrectangles representing the distribution uncertainty where our extended concepts of SHAP scores apply, and compare these scores against the usual, point-like SHAP score, in order to reveal cases where these new hypervolume scores are more informative than the traditional scheme. We expect our proposal to be able to detect features whose ranking is vulnerable to small distribution shifts, and aim to study how such sensitivity starts to vanish as we reduce the uncertainty over the distribution (i.e., as the hyperrectangles get smaller).
@@ 5.2 Methodology
We pick 20 different entities uniformly at random and compute the SHAP score for each of these entities and for each feature at all the vertices of the different hyperrectangles. Instead of using the polynomial-time algorithm for the SHAP value computation over decision trees, we actually compute an algebraic expression for the SHAP polynomial for each pair of entity and feature, simplify it, and then use it to compute any desired SHAP score given any product distribution.
@@ 5.3 Results
We can also observe in Figure 1 that something similar happens as well for feature housing_median_age. When $p=0.4\%$ its SHAP interval intersects with the intervals of the three features longitude, latitude and population. Nonetheless, by inspecting the difference of the scores at the vertices of $\mathcal{I}$ it can be seen that there is no point in which housing_median_age is ranked below the other features (i.e., housing_median_age dominates all three features). Meanwhile, if we compare two of the other features such as longitude and latitude, we observe that they have a relevant intersection even when $p=6.4\%$, and in 179 out of the 256 vertices latitude is ranked below longitude.
@@ 6 Conclusions
We have analyzed how SHAP scores for classification problems depend on the underlying distribution over the entity population when the distribution varies over a given uncertainty region. As a first and important step, we focused on product distributions and hyperrectangles as uncertainty regions, and obtained algorithmic and complexity results for fundamental problems related to how SHAP scores vary under these conditions. As a proof of concept, we showed through experimentation that considering uncertainty regions has an impact on feature rankings for a non-negligible percentage of the entities, and that the solutions to our proposed problems can provide insight on the relative rankings even in the presence of uncertainty.
Feature Independence.
We stress that from a complexity point of view, it is natural to start with product distributions. As shown in [27], the computation of SHAP scores becomes intractable for trivial classifiers as soon as we consider simple non-product distributions such as naive Bayes distributions. As a consequence, computing maxima of SHAP polynomials is trivially intractable in this setting. As we discussed at the beginning of Section 4.2, we focused on the cases for which computing SHAP scores is tractable, since this gives us the possibility to also obtain tractability for computing maxima of SHAP polynomials. We considered the prominent case of decomposable and deterministic Boolean circuits under product distributions. This case has been shown to be tractable in [27] and [3] (in [27] product distributions are referred to as fully-factorized distributions).
For our distributional shift analysis, considering more general distributions on the entity population would still require specifying a class of them and their regions of variation. A natural next step in this direction, that would build on our work, consists in imposing additional domain knowledge on the product distribution, leading, in particular, to certain dependencies among features. For example, a constraint specifying that “house lots located at the seaside have a price about \$2M". More generally, a conjunction $\varphi$ of such constraints is associated to an event $E^{\varphi}\subseteq{\tt ent}(X)$ containing all entities that satisfy it. Conditioning on this event leads to a new distribution $\mathbb{P}^{\prime}$ defined by $\mathbb{P}^{\prime}(A):=\mathbb{P}(A|E^{\varphi})$, for $A\subseteq{\tt ent}(X)$. The shift analysis would be done on the new entity space $\langle{\tt ent}(X),\mathbb{P}^{\prime}\rangle$, where (in general) features will not be fully independent anymore. This would be done, without having to start from scratch with $\mathbb{P}^{\prime}$, by taking advantage of our previous analysis of the original product distribution $\mathbb{P}$.
It is worth mentioning that imposing and exploiting domain semantics when defining and computing explanation scores is interesting in its own right (see [5] in relation to the RESP score). The problem of using domain knowledge in Explainable AI has been scarcely investigated in the literature.
Explanatory Robustness.
It would be interesting to know how a local perturbation of a given distribution in the uncertainty region affects the SHAP scores, or their rankings. This would be a proper robustness analysis with respect to the distribution (as opposed to how SHAP scores vary in a region). Of course, this is a different problem from the most common one of robustness with respect to the perturbation of an input entity (see e.g., [1, 13]).
There are several other problems left open by our work. The inclusion of non-binary features and labels is a natural next step. It would also be interesting to analyze others proposed scores (e.g., LIME [20], RESP [7], Kernel-SHAP [16]) in the setting of distributional uncertainty.
@@ Acknowledgments
This research has been supported by an STIC-AmSud program, project AMSUD210006. We appreciate the code made available by Jorge E. León (UAI, Chile) for SHAP computation with the same dataset used in our work.  Bertossi has been supported by NSERC-DG 2023-04650, the Millennium Institute for Foundational Research on Data (IMFD, Chile), and CENIA (Basal ANID FB210017, Chile).  Romero is funded by the National Center for Artificial Intelligence CENIA FB210017, Basal ANID. Pardal is funded by DFG VI 1045-1/1.  Abriola, Cifuentes and Pardal are funded by FONCyT, ANPCyT and CONICET, in the context of the projects PICT 01481 and PIP 11220200101408CO. Abriola is additionally funded by the project PIBAA 28720210100188CO.  Martinez is partially funded under the Spanish project PID2022-139835NB-C21 funded by MCIN/AEI/10.13039/501100011033, PIE 20235AT010 and iTrust (PCI2022-135010-2). Pardal was supported by the DFG grant VI 1045-1/1.
@@ Appendix A Appendix
Although the proof of this proposition can be found in [14], we show a different argument that we consider simpler.
@@
Theorem 3. Given a multilinear polynomial $f$, a hyperrectangle =[$\mathcal{I}={}_{i=1}^{n}[a_{i},b_{i}]$, and a rational $q$, the problem of deciding whether there is an $\textbf{x}\in\mathcal{I}$ such that $f(\textbf{x})\geq q$ is NP-complete.
@@
The problem is in NP, and the proof of this fact was already sketched in Section 4.1: due to Proposition 2, if $f$ attains a value larger than $q$, then there is a vertex x from $\mathcal{I}$ such that $f(\textbf{x})\geq q$. Therefore, a non-deterministic Turing Machine can guess which vertex it is and perform the evaluation.
@@
We prove the hardness of REGION-MAX-SHAP through a reduction from VERTEX-COVER. Let $G=(V,E)$ be a graph and $k\geq 1$ be an integer. We shall define a classifier $M$, an entity $e_{0}$, a feature $x_{0}$, an interval $\mathcal{I}$, and a bound $q$ such that $G$ has a vertex cover of size at most $k$ if and only if there is a $\textbf{p}\in\mathcal{I}$ such that $Shap_{M,e_{0},x_{0}}(\textbf{p})\geq q$.
The set of features is $X\vcentcolon=\{x_{0}\}\cup V\cup\{w\}$ where $V$ is the set of nodes of $G$ and $n\vcentcolon=|V|$. We consider the null entity $e_{0}$ that assigns 00 to all features, and the feature $x_{0}$ to define the SHAP polynomial. The hyperrectangle is $\mathcal{I}=[1,1]\times{}_{i=1}^{n}[0,1]\times[\varepsilon,\varepsilon]$. The value of $\varepsilon$ and the bound $q$ will be defined later. Recall that we can assume that the maximum of $Shap_{M,e_{0},x_{0}}$ is attained at a vertex of $\mathcal{I}$, i.e., at a point where $p_{x_{0}}=1$, $p_{w}=\varepsilon$ and $p_{v}\in\{0,1\}$ for all $v\in V$. The probabilities $(p_{v})_{v\in V}$ will be used to “mark” which nodes belong to the vertex cover. Formally, for a subset of nodes $C\subseteq V$ of $G$, we associate the vector $\textbf{p}^{C}=(p_{x})_{x\in X}$ such that $p_{x_{0}}=1$, $p_{w}=\varepsilon$, and for $v\in V$, we have $p_{v}=0$ if $v\in C$, and $p_{v}=1$ otherwise. Note that this gives us a bijection between subsets of nodes of $G$ and vertices of $\mathcal{I}$.
Intuitively, we devise the classifier $M$ such that $Shap_{M,e_{0},x_{0}}$ has the form
@@
where $\ell$ is the size of the subset $C$ and the term $T_{n,\ell}$ grows as $\ell$ grows. The terms $I_{u,v}$ can be made sufficiently large by choosing $\varepsilon$ properly. We choose the bound $q$ to be $T_{n,k}$. If $C$ is a vertex cover, then each term $p_{u}p_{v}I_{u,v}=0$ and hence the first sum of $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})$ is 00. On the other hand, if $C$ is not a vertex cover, we get a small negative value for $Shap_{M,e_{0},x_{0}}$, smaller than $q$ in particular, as we get penalized by at least one $I_{u,v}$ (from an uncovered edge $\{u,v\}$). Hence, the only way to obtain $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})\geq q$ is to pick $C$ to be, in the first place, a vertex cover, and secondly, one of size $\ell\leq k$.
@@
Recall $\ell$ denotes the size of $C$. We pick $\varepsilon\vcentcolon=\frac{c_{n-1}}{\sum_{i=0}^{k}\binom{k}{i}c_{n-i}}$ and $q\vcentcolon=-\varepsilon\sum_{i=0}^{k}\binom{k}{i}c_{n-i}=-c_{n-1}$. These values can be computed in $poly(n)$ because $k\leq n$. Similarly, the rest of the reduction can be performed in polynomial time.
Now we show the correctness of the reduction: $G$ has a vertex cover of size at most $k$ if and only if there is a $\textbf{p}\in\mathcal{I}$ such that $Shap_{M,e_{0},x_{0}}(\textbf{p})\geq q$. Suppose $G$ has a vertex cover of size at most $k$. Then it has a vertex cover $C$ of size precisely $k$. We can take the point $\textbf{p}^{C}$ since:
@@
$=q.$
Suppose now that there is no vertex cover of $G$ of size at most $k$. We claim that $Shap_{M,e_{0},x_{0}}(\textbf{p})<q$ for all $\textbf{p}\in\mathcal{I}$. It suffices to check this for the vertices of $\mathcal{I}$, that is, for all vectors of the form $\textbf{p}^{C}$ for subsets of nodes $C\subseteq V$. Now, take an arbitrary subset $C$. Suppose first that $C$ is not a vertex cover, and that the edge $\{u,v\}$ is not covered, then:
@@
Note that $I_{u,v}\geq c_{n-1}$ as $c_{n-1}$ participates in the sum when $S=(V\setminus\{u,v\})\cup\{w\}$. It follows that $Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})<-c_{n-1}=q$.
@@
$Shap_{M,e_{0},x_{0}}(\textbf{p}^{C})<-\varepsilon\sum_{i=0}^{k}\binom{k}{i}c_{n-i}=q.$
@@ A.3 The problem REGION-MIN-SHAP
Thus, $q$ is an upper bound for $Shap_{M,e,x}(\textbf{p})$ for $\textbf{p}\in\mathcal{I}$ if and only if $-q$ is a lower bound of $Shap_{1-M,e,x}(\textbf{p})$ over the same hyperrectangle. ∎
@@
The idea behind our construction is similar to the one in the proof of Theorem 5. In this case, we devise a model $M$ such that:
@@
Finally, let $\mathcal{D}(S)=\phi_{M^{\prime},e^{\prime}}(S\cup\{x\})-\phi_{M^{\prime},e^{\prime}}(S)=\mathcal{D}(S\setminus\{w\})$ and observe that:
@@
$=\sum_{\begin{subarray}{c}S\subseteq X^{\prime}\setminus\{x,w\}\end{subarray}}c_{|S|+1,n+1}\mathcal{D}(S)+c_{|S|,n+1}\mathcal{D}(S)$